OpenStudy (anonymous):

can somebody, anybody help? Find EF E Q D F P G 3x - 2 3x + 5 8x - 8 Reply Using Drawing Can you teach me how to do this?

OpenStudy (wolf1728):

What exactly is the question?

OpenStudy (anonymous):



OpenStudy (anonymous):

SORRY, Find EF

OpenStudy (anonymous):

CCan you teach me how to do this?

OpenStudy (jdoe0001):

hmm, I'm assuming QP is the Mid-Segment of the trapezoid? if so, midsegment of a trapezoid = \(\bf \cfrac{base1+base2}{2} \) you have 2 bases say base1 = 3x - 2 base2 = 8x - 8 and the midsegment = 3x + 5 thus => \(\bf \textit{midsegment } = \cfrac{base1+base2}{2} \implies 3x+5 = \cfrac{(3x-2)+(8x-8)}{2} \)

OpenStudy (anonymous):

I'm so confused, help!

OpenStudy (jdoe0001):

hmm, well.... I don't see anything confusing there :(

OpenStudy (jdoe0001):

we're assuming you know what a trapezoid is, and its midsegment is the line half-way to the top

OpenStudy (wolf1728):

Can you solve for x? 3x + 5 = (3x -2)/2 + (8x -8)/2

OpenStudy (anonymous):

yes I realize that, so 3x-2 falls right in between E&F so do I solve that for E & F? z\

OpenStudy (wolf1728):

I'd solve for x

OpenStudy (anonymous):

I've mever done 1 with 3 equations I've just done them with 1 on each side of the equal sign. Do you think you can help me work it out?

OpenStudy (wolf1728):

It is actually 1 equation and here's how to evaluate it: 3x + 5 = (3x-2)/2 + (8x -8)/2 How about getting out the fractions on the right hand side?

OpenStudy (anonymous):

3x+5=6x-4+16x-16, am I right so far?

OpenStudy (anonymous):

3x+5=22x+-20???

OpenStudy (wolf1728):

No - you multiplied by 2 instead of dividing. Here, I'll clear out 1 fraction for you: 3x + 5 = 1.5x -1 + (8x-8)/2 Can you clear out that right hand fraction? Remember to divide.

OpenStudy (anonymous):

3x+5=1.5x-1+4x-4 How are we looking?

OpenStudy (wolf1728):

Better No you would add all the terms together.

OpenStudy (anonymous):

so I have 7.5 I'm sorry 8.5x and 0=

OpenStudy (anonymous):

isn't any # divided by 0, 0?

OpenStudy (wolf1728):

No you got 8.5 because you added left side numbers with right side numbers. WAY back in my day they taught us to transpose the equation. So if you move 3x to the right side, it becomes -3x

OpenStudy (anonymous):

O I'm sorry I was moving everything to the right. please be patient I'll try it again

OpenStudy (wolf1728):

Remember whenever you move everything from one side to the other - change the signs.

OpenStudy (wolf1728):

Gotta head out. Be back in 30 minutes - maybe someone else can help you. You are close to finishing.

OpenStudy (anonymous):

Do I first add all the like terms on the right?

OpenStudy (anonymous):

O I'm sorry I was moving everything to the right. please be patient I'll try it again

OpenStudy (saifoo.khan):

Do you still need help?

OpenStudy (wolf1728):

Hey I'm back !!!

OpenStudy (anonymous):

Hello wolf

OpenStudy (saifoo.khan):

So where do you need help?

OpenStudy (anonymous):

3x+5=1.5x-1+4x-4 is this correct? Do I combine 1.5x & 4x then subtract -5.5x from +3x, leaving the sum on the left side of the equation? Then taking a negative 5 from the left side, adding the negatives giving me -10?

OpenStudy (anonymous):

10/2.5=4 is the correct answer to the equation x=4?